the probability that a patient recovers from a rare blood disease is 0.4. if 15 people are known to have this disease,what is the probability that at most 3 survive?
Answer by Boreal(15235)

That would be 0,1,2,3
probability 0 survive is 0.6^15=0.0005
probability 1 survives is 15C1 0.4*0.6^14=0.0047
probability 2 survive is 15C2*0.4^2*0.6^13=0.0219
probability 3 survive is 15C3*0.4^3*0.6^12=0.0634
That sum is 0.0905 ANSWER.